calculate the done when 50 grams of iron is dissolved in HCL at degree Celsius in ab open beaker when the atmospheric pressure is 1 atm

Grade:11

1 Answers

Arun
5 years ago
When iron reacts with dilute hydrochloric acid the iron dissolves and hydrogen is liberated. 

Fe + 2HCl = Fe Cl 2 + H2 ... (1) 

To answer the question you must consider how this process can result in work being done. An obvious answer is that if the reaction proceeds inside a closed cylinder fitted with a movable piston with atmospheric pressure acting on the outside, then the release of gas as a result of the reaction would move the piston outward against the atmospheric pressure. Work done could be calculated as V*Pa where V is the swept volume of the piston (equal to the volume of gas liberated), and Pa is the atmospheric pressure. This work would equally be done if the reaction occurred in an open beaker. If the piston were not allowed to move no work would be done; this would be the condition if the reaction occurs in a closed container of fixed volume. This latter point answers the first question; zero work done. 

To answer the second question we must calculate the volume of gas liberated by the stated reaction. 

According to (1), each mole of Fe produces 1 mole of H2. The number of moles of Fe involved in the reaction is 50/8.3145 = 6.0136, so this is the no of moles of H2. 

Each mole of an ideal gas has a volume of 24.465 L at a pressure of 1 standard atmosphere (101.3kPa), so the total volume of H2 is 24.465*6.0136 = 147.122L = 0.14712m^3 

The work done is then V*Pa = 0.14712*101.3 = 14.904kJ